OpenStudy (anonymous):

(x^(5/2) + 3/(sqrt(2) )^2

OpenStudy (anonymous):

And I find the first term as \[x^{5}\] and the last term as \[\frac{ 9 }{ 2 }\] but i don't understand how the middle term becomes \[3\sqrt{2}x ^{5/2}\]

OpenStudy (anonymous):

\[\left( x ^{5/2}+\frac{ 3 }{ \sqrt{2} } \right)^{2}\]

OpenStudy (anonymous):

\[\left( a+b \right)^2=a^2+b^2+2ab\] \[here~2*x ^{\frac{ 5 }{ 2 }}*\frac{ 3 }{ \sqrt{2} }=3\sqrt{2}x ^{\frac{ 5 }{ 2 }}\]

OpenStudy (triciaal):

This is the same just more of the missing steps

OpenStudy (anonymous):

\[2=\left( \sqrt{2} \right)^2\]

OpenStudy (anonymous):

i only solved 2ab

OpenStudy (anonymous):

when I multiply 2ab I am getting \[\frac{ 6\left( \sqrt{x} \right)^{5} }{ \sqrt{2} }\] why is this not right.? @surjithayer

OpenStudy (triciaal):